OpenStudy (gabylovesyou):

The sum of two consecutive even integers is 234. What is the larger integer? Explain how to do this pliss!

jimthompson5910 (jim_thompson5910):

let x = first integer, the next consecutive even integer is x+2 So x+x+2 = 234

OpenStudy (anonymous):

You need an equation, if x is the smaller integer and the larger integer is 2 bigger, how could you write the second intger using x?

OpenStudy (anonymous):

Let x = larger integer. This means the other number is x - 2. x - 2 + x = 234 2x - 2 = 234 2x = 236 x = 118

OpenStudy (anonymous):

Come on Snapbackalive, don't give them the whole answer, the point of this site is learning.

OpenStudy (gabylovesyou):

@jim_thompson5910 is it 116??

jimthompson5910 (jim_thompson5910):

x = 116 when you solve x+x+2 = 234 for x, very good

jimthompson5910 (jim_thompson5910):

so if x = 116, then what is x+2 ?

OpenStudy (gabylovesyou):

118

jimthompson5910 (jim_thompson5910):

good

OpenStudy (gabylovesyou):

Thank you!! :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

yes